OpenStudy (anonymous):

For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction

OpenStudy (kirbykirby):

A): 0*0=0 0*1 = 0 1*1 = 1 All the results of the operations are still in the set {0,1}, i.e. the set closed under multiplication B): 1/1=1 0/1 = 0 but 0/0 or 1/0 are not defined, so not closed C: 1+1 = 2, not in {0,1}: not closed D: 0-1=-1, not in {0,1} not closed

OpenStudy (kirbykirby):

essentially you need ALL elements under the operation to be still in the set {0,1} to be closed. If you find one pair of numbers in {0,1} under an operation which doesn't yield 0 or 1, then it's not closed under that operation

OpenStudy (anonymous):

@CrystalBurley

OpenStudy (anonymous):

its D subtraction